### v1

Abstract
In this paper, we introduce a novel discriminative loss function with large margin in the context of Deep Learning. This loss boosts the discriminative power of neural nets, represented by intra-class compactness and inter-class separability. On the one hand, the class compactness is ensured by close distance of samples of the same class to each other. On the other hand, the inter-class separability is boosted by a margin loss that ensures the minimum distance of each class to its closest boundary. All the terms in our loss have an explicit meaning, giving a direct view of the feature space obtained. We analyze mathematically the relation between compactness and margin term, giving a guideline about the impact of the hyper-parameters on the learned features. Moreover, we also analyze properties of the gradient of the loss with respect to the parameters of the neural net. Based on this, we design a strategy called partial momentum updating that enjoys simultaneously stability and consistency in training. Furthermore, we also investigate generalization errors to have better theoretical insights. Our loss function systematically boosts the test accuracy of models compared to the standard softmax loss in our experiments.
keywords:
Deep Learning, Loss Function, Large Margin Loss
1 Introduction
The standard approach to train a neural net for classification stands on a softmax loss. This loss consists of a softmax layer and the cross-entropy divergence. However, one of the main drawbacks of this loss is that it generally only helps the network to produce separable, but not sufficiently discriminative features ([1],[2]). In many problems, the intra-class variation is very large, meaning that the samples in each class are very diverse. But at the same time, the inter-class separability is small. That is, there exists some samples originating from different classes but they are very similar. This make the prediction much more difficult. Therefore, to achieve optimal generalization capability, a good machine learning algorithm in general, and a neural network in particular, should learn to produce features with high intra-class compactness and high inter-class separability [3].
To reach this goal, we introduce in this paper a novel loss function on the features of the penultimate layer (right before the softmax layer) in addition to the softmax loss. As this loss applies only on this penultimate layer, it can be used generically on any Neural Net model, for an end-to-end training, based on any gradient-based optimization method. Our new loss function is the combination of two terms: a hinged center loss and a margin loss. The hinged center loss ensures the compactness. By minimizing this loss, the resulted intra-class compactness of features is maximized. Notice that a center loss has been used in some previous works (e.g. [4]). This latter loss directly minimized the distance of each feature to its class centroid. This could encourage the model to a collapsed situation ([5]). That is, the model learns to project all the samples of each class to a sole point. To avoid this situation, one needs to add a quite complex penalty term and to be careful in the training process. On the contrary, in our method, inspired by the work of [6], we opt for hinged center loss, which only pushes the distance of each point to its centroid to be smaller than a predefined positive term $\delta_{v}$. This allows us to avoid the collapse phenomenon. On the other hand, the margin loss boosts the class margin. Here, the class margin of a given class is defined as its distance from the decision boundary. For this purpose, we first derive an exact analytical formula for the decision boundaries (see Section 4.1). Based on this, we also provide an analytical formula for the class margin and so we are able to maximize this margin. In addition, we derive some sharp mathematical insights on the relation between the hyperparameters both of the hinged center loss and the margin one, providing a lower bound for the class margin.
Ideal features should have their maximal intra-class distance smaller than their minimal inter-class distance ([2]), especially for classification problems where the samples in each class possess high variability while samples from different classes are quite similar. Enforcing the model to produce these ”ideal features”, it can learn more accurately the representative characteristics of each class (so that intra-class distance is small). At the same time, it focuses more on the characteristics that makes the difference between classes (so that the minimal inter-class distance is larger than the maximal intra-class distance). This boosts the discriminative power of the learned features. By deriving analytically both the relation between the compactness term and the margin term of our loss and the hyper-parameters, we can explicitly enforce the model to produce ideal features.
Besides, we directly compute the gradients of this loss with respect to the parameters of the neural net. This provides some insights about gradient. Some properties of the gradient leads to difficulty in updating the parameters. Based on this, we design a strategy, called partial momentum to overcome this drawback. This gives simultaneously, stability and consistency in the training process.
Our contributions can be summarized as follows:
  • The loss function provided here is the first one that enables both to model class compactness and margin simultaneously in a softmax model using an explicit formula (without any approximation).
  • We provide theoretical insights for a better understanding of feature learning in softmax models.
  • We conduct experiments on standard datasets. According to the quantitative results, our loss function systematically boosts the test accuracy compared to softmax loss, proving correctness of our insight. The qualitative results lead us to the same conclusion.
  • Thanks to our experiments, we find that by only boosting the discriminative power of the penultimate layer of a neural net, the features of the intermediate layers also become more discriminative.
2 Related works
The work [4] proposes a center loss applied in parallel to the standard softmax loss. This loss encourages the direct minimization of the distance between each point and the centroid of the corresponding class (in feature space). As discussed in [5], this could lead to feature collapse, where all the samples of each class collapse to a sole point. Therefore, one needs to add a penalty term and to be very careful in the training process to avoid this phenomenon. Here, we tackle the problem by using a hinged center loss, inspired by the work of De Brabandere et al. [6]. This only enforces the distance from centroid to be smaller than a certain predefined distance, avoiding feature collapse. Furthermore, center loss in [4] only explicitly encourages intra-class compactness and the inter-class margins are not taken into account. Hence, there is no guarantee about the margin.
The benefits of large-margin in the context of deep learning is pointed out in Liu et al. [7] and Liu et al. [2]. In these works the authors ignore the bias terms of the softmax layer, and consequently the margin can be viewed via angles between vectors. In contrast, our method makes no change in the softmax layer and we calculate explicitly the margin based directly on the Euclidean distance. Moreover, their method only encourages inter-class margin while the intra-class compactness is not explicitly considered. As discussed in the introduction, in scenarios where intra-variance class is large, intra-class compactness is necessary.
In the more recent work Zhou et al. [3], intra-class compactness and inter-class separability are both considered. However, this method completely ignores the magnitude of class prototypes to come up with the final loss. Moreover, this method only maximize the distance between class (in some sense), without considering the decision boundaries. It is clear that even with good inter-class separability, when the decision boundaries are not well adapted, for example too close to a class, good inter-class separability is no longer useful because a new example can easily cross to the other side of the decision boundary, leading to a poor classification.
In the work of Elsayed et al. [8], the class margin from the decision boundaries is considered. This method stands on the same ideas behind the classical SVM [9] in the sense that it imposes a minimum value for the margin of each class to the decision boundaries. Indeed, this paper proposes to boost the margins for the different layers of a neural net using a first-order approximation. In contrast, our method models the margin directly on the feature space of the penultimate layer. Our method therefore provides an exact formula for both the decision boundaries and the class margins from these boundaries. Moreover, the method proposed in Elsayed et al. [8] does not consider the intra-class compactness contrarily to ours.
Tang [10], also based on a margin approach, proposes to apply the multi-class SVM in the context of Deep Learning. However, this method is based on one-vs-rest approach, i.e. one needs to trained $C$ separate classifiers for $C$ classes. This makes the training much more heavy in the context of Deep Learning. In contrast, our method proposes a loss that consider all the classes simultaneously. Moreover, the intra-compactness is totally ignored therein.
3 Preliminaries and framework
Let us consider a classification problem with $C$ classes ($C\geq 2$). The input space is denoted by $\mathcal{X}$ (this space can be very complicated such as images, time series, vectors…). The neural net (backbone) transforms an input into a fixed-dimension vector. Formally we model the net by a function: $f_{\theta}:\mathcal{X}\mapsto\mathcal{F}\subseteq\mathbb{R}^{d}$, where $\theta$ is the set of parameters of the neural net, $d$ is the dimension of the so-called feature space $\mathcal{F}$. Given an input $x\in\mathcal{X}$, set $q=q(x)=f_{\theta}(x)$. In order to perform a classification task, $q$ is then passed through a softmax layer consisting of an affine (linear) transformation (Eq. (1)) and a softmax function (Eq. (2)):
$z=Wq+b,\ W\in\mathbb{R}^{C\times d}\ \text{and}\ b\in\mathbb{R}^{C},$ (1)
$\sigma(z)_{i}=\frac{e^{z_{i}}}{\sum_{j=1}^{C}e^{z_{j}}}.$ (2)
Here, for $i=1,\cdots,C,\;z_{i}$ is the $i^{th}$ component of column-vector $z$. Concatenating these two steps, we set $g(x)=\sigma(Wq+b)$. The predicted class is then the class with maximum value for $g$, i.e. $\hat{y}=\arg\max_{i}\sigma(z)_{i}$. Interestingly, notice that $\arg\max_{i}\sigma(z)_{i}=\arg\max_{i}z_{i}$. So that, $\hat{y}=\arg\max_{i}z_{i}$. In the standard approach, to train the neural net to predict maximal score for the right class, the softmax loss is used. This loss is written as
$\mathcal{L_{S}}=-\frac{1}{|\mathcal{B}|}\sum_{(x,y)\in\mathcal{B}}\log(g(f_{\theta}(x))_{y}).$ (3)
Here, $\mathcal{B}$ is the current mini-batch, $x$ being a training example associated with its ground-truth label $y\in\{1,2...,C\}$. By minimizing this loss function w.r.t. $\theta$ and $(W,b)$, the net learns to assign maximal score to the right class.
4 Large margin discriminative loss
4.1 Decision boundaries and the drawback of softmax loss
Consider the class pair $\{i,j\}$. We have:
$\begin{split}z_{i}-z_{j}&=(Wq+b)_{i}-(Wq+b)_{j}=\langle W_{i}-W_{j},q\rangle+(b_{i}-b_{j}).\end{split}$ (4)
Here, $z=\begin{pmatrix}z_{1}\\ \vdots\\ z_{C}\end{pmatrix}$ and $W=\begin{pmatrix}W_{1}^{T}\\ \vdots\\ W_{C}^{T}\end{pmatrix}$ with for $j=1,\cdots,C$, $W_{j}\in\mathbb{R}^{d}$. Set,
$\mathcal{P}_{ij}=\{q\in\mathcal{F},\langle W_{i}-W_{j},q\rangle+(b_{i}-b_{j})=0\}.$ (5)
[…]
4.2 Proposed Loss Function
To have a better classification, we are based on 2 factors: intra-class compactness and inter-class separability. To obtain these properties, we work in the feature space $\mathcal{F}$. In many classification problems, the intra-class variance is very large. So, by forcing the model to map various samples of the same class in a compact representation, the model learns the representative features of each class and ignores unhelpful details. Moreover, it may happen that samples in different classes are very similar. This leads to misclassification. Hence, we also aim to learn a representation having large margins between classes. In this way, the model learns the characteristics that make the difference between classes. Hence,it classifies better the samples. To achieve all these objectives, we propose the following loss function:
$\mathcal{L}=\alpha\cdot\mathcal{L}_{compact}+\beta\cdot\mathcal{L}_{margin}+\gamma\cdot\mathcal{L}_{reg}.$
Here, $\mathcal{L}_{compact}$, $\mathcal{L}_{margin}$ and $\mathcal{L}_{reg}$ force for class compactness, inter-class separability and regularization, respectively. We now discuss in detail these three terms. Let us consider the current mini-batch $\mathcal{B}$. Let $\mathcal{C}_{\mathcal{B}}$ be the set of classes in $\mathcal{B}$ and $\mathcal{C}_{c}^{\mathcal{B}}$ be the examples of class $c$ in $\mathcal{B}$. To ensure intra-class compactness, we use the discriminative loss proposed in De Brabandere et al. [6]. Note that in the latter article, this loss is used in the different context of image segmentation. This loss writes,
$\mathcal{L}_{compact}=\frac{1}{|\mathcal{C}_{\mathcal{B}}|}\times\sum_{c\in\mathcal{C}_{\mathcal{B}}}\frac{1}{|\mathcal{C}_{c}^{\mathcal{B}}|}\sum_{q\in\mathcal{C}_{c}^{\mathcal{B}}}[||m_{c}-q||-\delta_{v}]_{+}^{2}.$ (6)
Here, ||.||||.||| | . | | is the $L2$ distance, $[q]_{+}=\max(0,q)$ and $m_{c}$ is the centroid of the class $c$. This function is zero when $||m_{c}-q||<\delta_{v}$. Hence, this function enforces that the distance of each point to its centroid is smaller than $\delta_{v}$. Notice that this function only pushes the distance to be smaller than $\delta_{v}$ and not to be zero. Hence, we avoid the phenomenon of mode collapse.
To have a better inter-class separability, we build a loss function enforcing large margin between classes and at the same time taking into account the decision boundaries. A naive strategy would be to maximize distance of each sample to all the decision boundaries (see for example Elsayed et al. [8]). However, this is very costly and not really necessary. Instead, we propose to maximize the distance of each centroids to the decision boundaries. Indeed, we will give in Proposition 2 a lower bound for the class margins. The margin loss is defined as follows,
$\mathcal{L}_{margin}=\frac{1}{|\mathcal{C}_{\mathcal{B}}|}\times\sum_{c\in\mathcal{C}_{\mathcal{B}}}\max_{i\neq c}\left([\delta_{d}+d(m_{c},\mathcal{P}_{ci})\text{ sign}(g(m_{c})_{i}-g(m_{c})_{c})]_{+}\right).$ (7)
[…]
$d(m_{c},\mathcal{P}_{ci})=\frac{|\langle W_{c}-W_{i},m_{c}\rangle+(b_{c}-b_{i})|}{||W_{c}-W_{i}||}.$ (8)
Our loss function encourages each centroids to be far away from the decision boundaries. However, there are no guarantee that the decision boundaries lead to closed cells. The resulted centroids could be pushed far away. Hence, to address this problem, we add a regularization term as proposed in De Brabandere et al. [6],
$\mathcal{L}_{reg}=\frac{1}{|\mathcal{C}_{\mathcal{B}}|}\sum_{c\in\mathcal{C}_{\mathcal{B}}}||m_{c}||.$ (9)
4.3 Properties of intra-class compactness and inter-class separability
In this section, we investigate the properties of compactness and separability of the loss function. Furthermore, we discuss the impact of the hyper-parameters $\delta_{v}$ and $\delta_{d}$. This gives us a guideline on the choice of these hyper-parameters.
Definition 4.1 (Class dispersion).
Let us define the dispersion of a given class $c$ as the maximal distance between two samples in this class: $\operatorname{dispersion}(c)=\max_{p,q\in\mathcal{C}_{c}}d(p,q).$
Proposition 1.
If $\mathcal{L}_{compact}=0$, then the dispersion of all classes is at most $2\delta_{v}.$
Proof.
See Appendix A.1. ∎
This last proposition shows that the hinged center loss ensures the intra-compactness property of each class.
Definition 4.2 (Class margin).
Let us define the margin of a given class $c$ as the smallest distance of samples in this class to its closest decision boundary, i.e.
$\operatorname{margin}(c)=\min_{q\in\mathcal{C}_{c}}\left(\min_{i\neq c}d(q,\mathcal{P}_{ci})\right)$
Proposition 2.
Assume that $\mathcal{L}_{compact}=\mathcal{L}_{margin}=0$. Then,
  1. 1.
    If $\delta_{d}>\delta_{v}$, then the margin of all classes is at least $\delta_{d}-\delta_{v}$.
  2. 2.
    If $\delta_{d}>2\delta_{v}$, then the distances between any points in the same class are smaller than the distances between any points from different classes.
Proof.
See Appendix A.2. ∎
Hence, if we aim to obtain class margin at least $\varepsilon$, then we can set $\delta_{d}=\delta_{v}+\varepsilon$. Furthermore, this proposition provides a guideline for the choice of $\delta_{v}$ and $\delta_{d}$. We are aiming for a representation with not only a large inter-class margin, but also one in which the distances between points in the same class are smaller than the distances between points in different classes. This is particularly useful for problems where samples in each class are too diverse whereas samples from different classes are too similar. Indeed, by enforcing this property in the feature space, the model learns more useful representative characteristics of each class. So that, intra-class distance is small even for dissimilar samples of the same class. At the same time, the model focuses more on the characteristics that makes the difference between classes. In this way, even very similar samples but coming from different classes are better separated.
4.4 Partial momentum for centroids
Let us now consider a class $c$. To compute the centroid of this class, there are 2 straightforward ways:
  • Naive way. using all the sample of the considered class in the current mini-batch: $m^{t}_{c}:=m^{current}_{c}=\frac{1}{|\mathcal{C}_{c}^{\mathcal{B}}|}\sum_{q\in\mathcal{C}_{c}^{\mathcal{B}}}q$.
  • Using momentum. $m^{t}_{c}:=m^{momentum}_{c}=\gamma\cdot m^{t-1}_{c}+(1-\gamma)\cdot m^{current}_{c}$, where $\gamma$ is chosen to be very close to $1$, such as $0.99$.
One major advantage of using momentum is stability. In fact, as we work with mini-batches, it can happen that the centroid of each class moves too much from one batch to another. In such case, we do not have a stable direction to that centroid. As $\mathcal{L}_{compact}$ aims to push each point to its corresponding centroid, the optimization becomes less effective. Thus, the use of momentum allows us to avoid this problem. However, using momentum makes the gradient much smaller when updating the model parameters. More precisely, we have following proposition:
Proposition 3.
$\nabla_{\theta}\mathcal{L}_{margin}^{moment}=(1-\gamma)\cdot\nabla_{\theta}\mathcal{L}_{margin}^{naive}$. Here, $\mathcal{L}_{margin}^{naive}$ and $\mathcal{L}_{margin}^{moment}$ are computed using the centroids updated based on naive way and momentum way, respectively.
Proof.
See Appendix B. ∎
This proposition shows that using centroid with or without momentum gives the same gradient direction. Nevertheless, with momentum the very small shrinking scaling factor $1-\gamma$ appears.
Further, this small gradient is multiplied by a small learning rate ( typically in the range $[10^{-5},10^{-2}]$). So, on the one hand, the parameter updating in the momentum method is extremely small (or even get completely canceled out by the computer rounding limit or machine epsilon). On the other hand, as discussed previously, using momentum allows more stability. To overcome the gradient drawback but to conserve the stability benefit, we combine the naive and momentum ways. We come up with a strategy named partial momentum. This strategy uses momentum for the compactness loss and naive way for the margin loss, respectively. Doing so, we have stable centroids. So that, each point is pushed in a stable direction. But at the same time, the centroids are kept consistent. That is, the parameters of the neural net evolve along training with sufficiently large gradients.
4.5 Squared loss or not?
We can notice that each term under the sum operation in $\mathcal{L}_{compact}$ is squared, whereas this is not the case for $\mathcal{L}_{margin}$. Indeed, squared of each term leads to more relaxing loss than non-squared version. More formally, when the term under the sum operation of $\mathcal{L}_{compact}$ and $\mathcal{L}_{margin}$ is still activated, its general form can be written as
$f(u)=\begin{cases}\pm||u-u_{ref}||+b,&\text{if not squared }\\ (\pm||u-u_{ref}||+b)^{2},&\text{if squared }\end{cases}$
where $u_{ref}$ is the reference point. In the case of $\mathcal{L}_{compact}$, $||u-u_{ref}||$ is the distance from a generic point to its corresponding centroid denoted here by $u_{ref}$. In the case of $\mathcal{L}_{margin}$, $||u-u_{ref}||$ is the distance of a centroid to its projection on the closest boundary. For sake of simplicity, we ignore here the sign before $||u-u_{ref}||$ as here this does not matter. So, we have:
$\nabla_{u}f=\begin{cases}\frac{u-u_{ref}}{||u-u_{ref}||},&\text{non squared case }\\ 2(||u-u_{ref}||+b)\times\frac{u-u_{ref}}{||u-u_{ref}||},&\text{squared case}\end{cases}.$
$\implies||\nabla_{u}f||=\begin{cases}1,&\text{not squared }\\ 2\times\bigg{|}||u-u_{ref}||+b\bigg{|},&\text{squared }\end{cases}.$
That is, if not squared, the gradient remains with a constant magnitude as long as it is still activated. In contrast, if the term is squared, then, for states close to be deactivated, $\left|||u-u_{ref}||+b\right|$ is close to 0. Hence, the gradient is very small. Thus, the magnitude of the update direction for $u$ becomes minimal when it is close to the deactivated state. Thus, on the one hand, squaring in $\mathcal{L}_{compact}$ makes it more relaxing. For example, if there exist some points too abnormal, then this condition does not enforce completely the point to be in the hyper-sphere around its centroid. On the other hand, by not squaring for the margin loss, we enforce harder the centroid until it attains the deactivated state. That is, its distance to the closest boundary is at least larger than $\delta_{d}$. This is important as the position of each centroid impact the distribution of the whole class. This insight justifies our proposed loss functions.
4.6 Generalization error
In this section, we investigate the generalization error of our method. For this, we first recall the notion of margin loss introduced in [11].
Definition 4.3 (Margin loss function).
For any $\rho>0$, the $\rho$-margin loss is the function $L_{\rho}:\mathbb{R}\times\mathbb{R}\mapsto\mathbb{R}_{+}$ defined for all $y,y^{\prime}\in\mathbb{R}$ by $L_{\rho}(y,y^{\prime})=\Phi_{\rho}(yy^{\prime})$ with,
$\Phi_{\rho}(x)=\begin{cases}1&\text{if}\ x\leq 0\ .\\ 1-x/\rho&\text{if}\ 0\leq x\leq\rho\ .\\ 0&\text{if}\ \rho\leq x\ .\end{cases}$ (10)
Interpretation. Considering $y\in\{-1,+1\}$ as ground-truth label and $y^{\prime}:=h(x)$ as the decision function, correct predictions correspond to the points with $yh(x)>0$. Therefore, this loss penalizes the points with wrong decision (i.e. $yh(x)<0$) or with correct decision but with confidence level $yh(x)<\rho$. Hence, $\rho$ can be regarded as a confidence parameter. If we have no demand about confidence level in the decision, we can set $\rho=0$. In this case, the margin loss becomes the function $1_{\{h(x)y<0\}}$, which penalizes the wrong predictions ($yh(x)<0$).
We will set our main theoretical results. The next subsections give and discuss non asymptotic bound theoretical risks.
4.6.1 Pairwise classification generalization error
We note that multi-classification can be seen as separate classifications in pairs. In other words, each class has to be separated from the other classes. Therefore, studying pairwise classification helps us to better understand the problem. For this end, we assume here that the input $(X,Y)$ ($X\in\mathcal{X}$ and $Y\in\{-1,1\}$), consists in a binary mixture and we aim to quantify the binary classification error of our method. The next theorem relates this error to the empirical one obtained on the training sample. Let $S$ denote the sample consisting in $N>0$ independent copies of $(X,Y)$. Recall that $\mathcal{F}$ denotes the feature space and let $\mathcal{M}(\mathcal{X},\mathcal{F})$ be the set of all measurable functions from $\mathcal{X}$ to $\mathcal{F}$. Let $R>0$ and $m_{1},m_{2}\in\mathcal{F}$ such that $\|m_{i}\|\leq R,\;(i=1,2)$. Given $r>0$, our oracle bounds involve the following functional set
$\mathcal{G}_{1}=\left\{f\in\mathcal{M}(\mathcal{X},\mathcal{F}):\sup_{x\in\mathcal{X}}\min\left(\|f(x)-m_{1}\|,\|f(x)-m_{2}\|\right)\leq r\right\}.$
Our first theorem writes,
Theorem 1.
For $h\in\mathcal{M}(\mathcal{X},\mathcal{F})$, let $R(h)=\mathbb{E}[1_{\{Yh(X)<0\}}]$ and $\widehat{R}_{S,\rho}(h)=\frac{1}{N}\sum_{i=1}^{N}\Phi_{\rho}(y_{i}h(x_{i}))$. Then,
  1. 1.
    Given a fixed mapping $f\in\mathcal{G}_{1}$ and $\Gamma>0$, let $H_{f}=\left\{\langle f(\cdot),w\rangle+b:\left\|\begin{pmatrix}w\\ b\end{pmatrix}\right\|\leq\Gamma\right\}$. For any $\delta>0$, with probability at least $1-\delta$, the following inequality holds,
    $R(h)\leq\widehat{R}_{S,\rho}(h)+\frac{2\Gamma\sqrt{(r+R)^{2}+1}}{\rho\sqrt{N}}+3\sqrt{\frac{\log\frac{2}{\delta}}{2N}},\;\;(h\in H_{f}).$ (11)
  2. 2.
    Let $H_{1}=\bigcup_{f\in\mathcal{G}_{1}}H_{f}$. For any $\delta>0$, with probability at least $1-\delta$ we have,
    $R(h)\leq\widehat{R}_{S,\rho}(h)+\frac{2\Gamma\sqrt{(r+R)^{2}+1}}{\rho}+3\sqrt{\frac{\log\frac{2}{\delta}}{2N}},\;\;(h\in H_{1}).$ (12)
Proof.
See Appendix C.1. ∎
Remarks. First of all, it is natural to work with the functional class $\mathcal{G}_{1}$, since the loss function (8) imposes the mapping of the input to the balls of each class in the feature space. Secondly, assuming that $\|m_{1}\|$ and $\|m_{2}\|$ are both bounded by $R$ is a reasonable assumption in view of the regularization term (9). In addition, to bound the theoretical loss, we use the empirical margin loss instead of our own. However, it is obvious that minimizing our loss leads to minimize the empirical margin loss. In fact, using our loss tends to separate the two classes and increase margins as shown in Proposition 2. Lastly, notice the difference between the two upper bounds provided in the last theorem. The first is local because $f$ is fixed, whereas the second is valid for all $f$ This explains the degradation in the second term of Eq. (12).
4.6.2 Generalization error of mapping each point to a hyper-sphere
In the above subsection, we have considered the pairwise classification. Furthermore, in our method, the loss contains a compactness component that enforces each point to be mapped in a hypersphere centered on the centroid of its class (in feature space). By imposing the model to satisfy this constraint on the training set, we expect to have the same property on the test set. Is this a reasonable objective? To answer this question, we now fix a particular class and quantify the mapping error. That is, the probability that a point is projected outside the correct hypersphere. The following theorem provides an empirical upper bound for this probability. Here, $S$ denote the sample consisting in $N>0$ independent copies of $X$. Let $R,r,\Lambda>0$, our oracle bound involves the following functional set
$H_{2}=\left\{r^{2}-||f(\cdot)-m||^{2}:\;||m||\leq R,f\in\mathcal{M}(\mathcal{X},\mathcal{F}),\;\sup_{x\in\mathcal{X}}\|f(x)\|\leq\Lambda\right\}.$
Our second theorem writes,
Theorem 2.
For any $\delta>0$, with probability at least $1-\delta$ over the draw of an i.i.d. sample $S$ of size $N$, for any $h\in H_{2}$, we have,
$\mathbb{P}(h(X)<0)\leq\widehat{R}^{\prime}_{S,\rho}(h)+\frac{2}{\rho}(\Lambda^{2}+2R\Lambda+\frac{R^{2}}{\sqrt{N}})+3\sqrt{\frac{\log\frac{2}{\delta}}{2N}}\ .$ (13)
Here, $\widehat{R}^{\prime}_{S,\rho}(h)=\frac{1}{N}\sum_{i=1}^{N}\Phi_{\rho}(h(x_{i}))$ is the empirical margin loss on $S$.
Proof.
See Appendix C.2. ∎
Remarks. Notice that if an input $x$ is mapped into a point inside the hyper-sphere (in the feature space), then $h(x)>0$. Hence, $\mathbb{P}(h(X)<0)$ measures the average error, i.e. examples whose mapping lies outside the hyper-sphere.
Notice also that on one hand, when $\rho=0$, $\widehat{R}^{\prime}_{S,\rho}(h)$ only penalizes training examples whose mapping is outside the hyper-sphere ($h(x)<0$). On the other hand, when $\rho>0$, it additionally penalizes examples mapped inside the hyper-sphere but having a margin from the hyper-sphere boundary less than $\rho$. We also remark that the upper bound (13) gets smaller as the number of training examples $N$ gets larger. Obviously, this is expected, as with larger $N$, the training examples cover better the underlying input distribution. Hence, the model tends to well behave on the test set, if it is well trained (i.e. small empirical loss on training set).
5 Experiments
In this section, we perform experiments on 2 standard datasets: CIFAR10 [12] and The Street View House Numbers (SVHN) [13]. CIFAR10 contains 50000 training images and 10000 test images of 10 classes (including vehicles and animals). SVHN has 73257 training images and 26032 for testing. We use ResNet18 [14] as backbone, followed by some fully connected layers prior to softmax layer.
5.1 CIFAR10
First, we qualitatively evaluate how our loss function helps to boost class compactness and inter-class separability compared to the softmax loss. To this end, we first train a model on the training set of CIFAR10. Next, we compute the features of the test set. Finally, we use the technique of t-SNE [15] for visualizing feature in a 2D space. The results are shown in Fig. 1(a) and 1(b). Each color represents a class. We see that by using softmax loss (Fig. 1(a)), the classes are not so well separated in the feature space. On the contrary, our method allows better inter-class separation (Fig.1(b)).
(a) Softmax
(b) Ours
Figure 1: t-SNE of CIFAR10 test images in the feature space after training with softmax loss and our loss function.
Next, we evaluate how our loss function helps the model generalize better on the test set by comparing test accuracies. We use only a sample fraction of the training set for training. Then we evaluate the model on the full test set, comparing softmax loss to ours. The results are shown in Table 1 and Fig. 2(a). We find that at different percentages of the training set, our method consistently performs significantly better than softmax loss in terms of test accuracy. Moreover, with only $60\%$ of the learning set, our methods already outperform the softmax loss, which uses the full training set. This result proves that our loss function helps the model to have a better generalization ability with less training data.
Table 1: Accuracy on test set of CIFAR10 at different training percentage of training set.
[…]
softmax 69.0 78.8 87.0 88.5 89.4 91.3
ours 72.2 82.8 91.0 91.9 92.4 93.3
(a) CIFAR10
(b) SVHN
Figure 2: Test accuracy of CIFAR10 and SVHN after training using only a certain fraction of training set with softmax loss and our loss.
5.2 SVHN
We perform the same experiments as the one done for CIFAR10. From Fig. 3(a) and 3(b), we see that our method helps to learn features with better separation and compactness. This once again shows correctness of our insights.
(a) Softmax
(b) Ours
Figure 3: t-SNE of SVHN test images in the feature space after training with softmax loss and our loss function.
The test accuracy using different percentages of training data is given in Table 2 and Fig. 2(b). Once again, our method systematically outperforms softmax loss. Remarkably, with only $40\%$ of the training set, our method already gives the same test accuracy as the softmax model using the full training set.
Table 2: Accuracy on test set of SVHN at different training percentage of training set.
[…]
softmax 86.9 90.7 93.2 93.8 94.2 95.1
ours 90.9 93.4 95.1 95.7 96.0 96.3
5.3 Impact of our loss function on intermediate layers
In the two experiment above, we use ResNet18 [14] as backbone, followed by some fully connected layers prior to softmax layer. Recall that our loss function is applied on the feature of the penultimate layer (right before softmax layer). Hence, it is interesting to see the impact of this loss function on the model. Indeed, the results shown in Figures 1 and 3 are the features of penultimate layer. It could be argued that our loss function only impacts the last fully connected layers to provide discriminating features, and does not really impact the backbone. To answer this question, we perform the same t-SNE technique for the intermediate layers of the backbone. ResNet18 includes 4 main blocs, each gives outputs of dimension $H_{i}\times W_{i}\times C_{i},\;(i=1,\cdots,4)$. $H_{i}$ and $W_{i}$ are spatial dimensions (and so depend on the input dimension). $C_{i}$ is the number of channels of the bloc $i$ (independent of the input dimension). For each input, we first perform a Global Average Pooling over the spatial dimensions to obtain a feature vector of dimension $C_{i}$ (for each bloc $i$). Then t-SNE is performed as before. The results for CIFAR10 and SVHN are shown in Figures 4 and 5.
(a) Bloc $1$
(b) Bloc $2$
(c) Bloc $3$
(d) Bloc $4$
(e) Bloc $1$
(f) Bloc $2$
(g) Bloc $3$
(h) Bloc $4$
Figure 4: t-SNE of CIFAR10 test images for different intermediate layers with softmax loss and our loss function. Top row: softmax loss. Bottom row: our loss.
(a) Bloc $1$
(b) Bloc $2$
(c) Bloc $3$
(d) Bloc $4$
(e) Bloc $1$
(f) Bloc $2$
(g) Bloc $3$
(h) Bloc $4$
Figure 5: t-SNE of SVHN test images for different intermediate layers with softmax loss and our loss function. Top row: softmax loss. Bottom row: our loss.
For the two datasets, we observe that there is no distinctive clusters in the first two blocs. This is expected as these are shallow layers, and no semantic feature is really captured. However, from the bloc 3, our loss function seems to help the model to produces more distinctive clusters compared to softmax loss (Fig. 4(g) vs Fig. 4(c) for CIFAR10 and Fig. 5(g) vs Fig. 5(c) for SVHN). This effect is well observed for the bloc 4 for both datasets (Fig. 4(h) vs Fig. 4(d) for CIFAR10 and Fig. 5(h) vs Fig. 5(d) for SVHN). This implies that our loss function not only impacts the feature of the penultimate layer, but really helps the model to learn more discriminative features at the intermediate levels. In conclusion, our loss function is powerful in the classification task by neural nets.
6 Conclusion
In this paper, we introduce a loss function for classification problems using softmax models. This loss is applied directly in the feature space of the penultimate layer. Hence, it can be used generically to boost the compactness and inter-class margins for better classification. We also give insights on our loss function for better understanding our method. The qualitative evaluations using t-SNE clearly show that our loss encourages more discriminative features. This is not only the case for the penultimate layer, but also for the intermediate layers. The numerical results once again confirm efficacy of our method. Indeed, using only a small part of training set with our loss function already gives the same test accuracy as training on the full training set with the softmax loss. In conclusion, the loss proposed here can be a plug-and-play tool to improve the performance of any classification task using softmax models.
Supplementary Material
Appendix A Proofs of intra-class compactness and inter-class separability
A.1 Proof of Proposition 1
Proof.
Let $q_{1}$ and $q_{2}$ be 2 arbitrary points in any class $c$. By triangle inequality, we have: $||q_{1}-q_{2}||\leq||q_{1}-m_{c}||+||m_{c}-q_{2}||\leq\delta_{v}+\delta_{v}=2\delta_{v}.$
A.2 Proof of Proposition 2
Proof of Proposition 2.1
Proof.
Consider an arbitrary class $c$. Let $p$ be a point in this class and let $q$ be any point on the hyperplane $\mathcal{P}_{ci}$ for any $i\neq c$. Then, by triangle inequality, we have:
$||m_{c}-q||\leq||m_{c}-p||+||p-q||.$
As $d(m_{c},\mathcal{P}_{ci})\geq\delta_{d}$, we have $||m_{c}-q||\geq\delta_{d},\ \forall q\in\mathcal{P}_{ci}$. At the same time, we also have $||m_{c}-p||\leq\delta_{v}$. Consequently, we obtain:
$\delta_{d}\leq||m_{c}-q||\leq||m_{c}-p||+||p-q||\leq\delta_{v}+||p-q||.$
Hence, $||p-q||\geq\delta_{d}-\delta_{v},\ \forall y\in\mathcal{P}_{ci}$. So, by definition, $d(p,\mathcal{P}_{ci})\geq\delta_{d}-\delta_{v}$. This holds $\forall p\in\mathcal{C}_{c}$ and $\forall i\neq c$. Hence, $\text{margin}(c)=\min_{p\in\mathcal{C}_{c}}\left(\min_{i\neq c}d(p,\mathcal{P}_{ci})\right)\geq\delta_{d}-\delta_{v}.$ As we choose an arbitrary class $c$, this holds for all classes. The proposition is proved.
Proof of Proposition 2.2
Proof.
Let $q_{1}$ and $q_{2}$ be 2 arbitrary points in any class same $c$. By Proposition 1, we have: $||q_{1}-q_{2}||\leq 2\delta_{v}.$ Now, let $p$ and $q$ by 2 arbitrary points in any two different classes $i$ and $j$, respectively. It suffices to show that $||p-q||>2\delta_{v}$ with $\delta_{d}>2\delta_{v}$. Again, by triangle inequality, we have:
$||m_{i}-m_{j}||\leq||m_{i}-p||+||p-m_{j}||\leq||m_{i}-p||+(||p-q||+||q-m_{j}||).$
So, $||p-q||\geq||m_{i}-m_{j}||-(||m_{i}-p||+||q-m_{j}||)\geq||m_{i}-m_{j}||-2\delta_{v}.$ Now, let us consider $||m_{i}-m_{j}||$. By Cauchy-Schwart inequality, we have:
$\left\lvert\langle m_{i}-m_{j},\frac{W_{i}-W_{j}}{||W_{i}-W_{j}||}\rangle\right\rvert\ \leq||m_{i}-m_{j}||\cdot\frac{||W_{i}-W_{j}||}{||W_{i}-W_{j}||}=||m_{i}-m_{j}||.$
Hence,
$\begin{split}||m_{i}-m_{j}||&\geq\left\lvert\frac{\langle W_{i}-W_{j},m_{i}\rangle}{||W_{i}-W_{j}||}-\frac{\langle W_{i}-W_{j},m_{j}\rangle}{||W_{i}-W_{j}||}\right\rvert\\ &=\left\lvert\frac{\langle W_{i}-W_{j},m_{i}\rangle+(b_{i}-b_{j})}{||W_{i}-W_{j}||}-\frac{\langle W_{i}-W_{j},m_{j}\rangle+(b_{i}-b_{j})}{||W_{i}-W_{j}||}\right\rvert\end{split}$
As $m_{i}$ and $m_{j}$ are on 2 different sides of the hyperplane $\mathcal{P}_{ij}$ (which is the decision boundary), $\langle W_{i}-W_{j},m_{i}\rangle+(b_{i}-b_{j})$ is of opposite sign of $\langle W_{i}-W_{j},m_{j}\rangle+(b_{i}-b_{j})$. Hence,
$\begin{split}\left\lvert\frac{\langle W_{i}-W_{j},m_{i}\rangle+(b_{i}-b_{j})}{||W_{i}-W_{j}||}-\frac{\langle W_{i}-W_{j},m_{j}\rangle+(b_{i}-b_{j})}{||W_{i}-W_{j}||}\right\rvert&=\left\lvert\frac{\langle W_{i}-W_{j},m_{i}\rangle+(b_{i}-b_{j})}{||W_{i}-W_{j}||}\right\rvert+\\ &\left\lvert\frac{\langle W_{i}-W_{j},m_{j}\rangle+(b_{i}-b_{j})}{||W_{i}-W_{j}||}\right\rvert.\end{split}$
Consequently,
$\begin{split}||m_{i}-m_{j}||&\geq\left\lvert\frac{\langle W_{i}-W_{j},m_{i}\rangle+(b_{i}-b_{j})}{||W_{i}-W_{j}||}\right\rvert+\left\lvert\frac{\langle W_{i}-W_{j},m_{j}\rangle+(b_{i}-b_{j})}{||W_{i}-W_{j}||}\right\rvert\\ &=d(m_{i},\mathcal{P}_{ij})+d(m_{j},\mathcal{P}_{ij})\geq 2\delta_{d}.\end{split}$
So, $||p-q||\geq||m_{j}-m_{j}||-2\delta_{v}\geq 2\delta_{d}-2\delta_{v}$. With $\delta_{d}>2\delta_{v}$, we have $||p-q||>2\delta_{v}$. So, $||p-q||>||q_{1}-q_{2}||.$
Appendix B Proof of Gradients in Proposition 3
Proof.
If $\mathcal{L}_{margin}=0$, then the gradient w.r.t. $\theta$ is 0, hence proposition is proved. Let us now consider the case where $\mathcal{L}_{margin}>0$ is still larger than 00. We can easily see that
$\nabla_{\theta}\mathcal{L}_{margin}=\text{ sign}(g(m_{c})_{i}-g(m_{c})_{c})\cdot\nabla_{\theta}d(m_{c},\mathcal{P}_{cj}).$
Hence, if we show $\nabla_{\theta}d(m_{c}^{current},\mathcal{P}_{ci})=(1-\gamma)\cdot\nabla_{\theta}d(m^{moment}_{c},\mathcal{P}_{ci})$, the proposition is proved. First, we recall that for two differentiable functions: $g_{1}:\mathbb{R}^{d_{1}}\mapsto\mathbb{R}^{d_{2}}$ and $g_{2}:\mathbb{R}^{d_{2}}\mapsto\mathbb{R}^{d_{3}}$, and if $g$ denotes the composed function $g:=g_{2}\circ g_{1}$, then for $x\in\mathbb{R}^{d_{1}}$ we have the following chain rule for computing the Jacobian matrix:
$J_{g}(x)=J_{g_{2}}(g_{1}(x))\times J_{g_{1}}(x).$
In the case where $d_{3}=1$, then we have:
$\nabla_{x}g=J_{g}(x)^{T}=J_{g_{1}}^{T}(x)\times J_{g_{2}}^{T}(g_{1}(x))=J_{g_{1}}^{T}(x)\times\nabla_{g_{1}(x)}g_{2}.$ (14)
Now, recall that $d(m_{c},\mathcal{P}_{ci})=\frac{|\langle W_{c}-W_{i},m_{c}\rangle+(b_{c}-b_{i})|}{||W_{c}-W_{i}||}$. For sake of brevity, we consider the case where $\langle W_{c}-W_{i},m_{c}\rangle+(b_{c}-b_{i})\geq 0$ and the argument is the same for the case $\langle W_{c}-W_{i},m_{c}\rangle+(b_{c}-b_{i})<0$. We have, $d(m_{c},\mathcal{P}_{ci})=\frac{\langle W_{c}-W_{i},m_{c}\rangle+(b_{c}-b_{i})}{||W_{c}-W_{i}||}.$ Using the chain rule from Eq. (14) with $g_{2}(\cdot)=\frac{|\langle W_{c}-W_{i},\cdot\rangle+(b_{c}-b_{i})|}{||W_{c}-W_{i}||}$, we get:
$\nabla_{\theta}d(m_{c}^{current},\mathcal{P}_{ci})=J^{T}_{m_{c}}(\theta)\times\nabla_{m_{c}}g_{2}.$
On the one hand, we can easily show that $\nabla_{m_{c}}g_{2}$ is independent of $m_{c}$ and equals to $\frac{W_{c}-W_{i}}{||W_{c}-W_{i}||}$. So,
$\nabla_{\theta}d(m_{c},\mathcal{P}_{ci})=J^{T}_{m_{c}}(\theta)\times\frac{W_{c}-W_{i}}{||W_{c}-W_{i}||}.$
On the other hand, we have that $m^{moment}_{c}=\gamma\cdot m^{t-1}_{i}+(1-\gamma)\cdot m^{current}_{c}$. Consequently,
$J^{T}_{m_{c}^{moment}}(\theta)=(1-\gamma)\cdot J^{T}_{m_{c}^{current}}(\theta).$
Hence, $\nabla_{\theta}d(m_{c}^{moment},\mathcal{P}_{ci})=(1-\gamma)\cdot\nabla_{\theta}d(m_{c}^{current},\mathcal{P}_{ci})$. So,
$\nabla_{\theta}\mathcal{L}_{margin}^{moment}=(1-\gamma)\cdot\nabla_{\theta}\mathcal{L}_{margin}^{naive}.$
Appendix C Proofs of generalization errors
In this section, we provide the complete proofs concerning the generalization errors discussed in Section 4.6. Our proofs are inspired by the methodology developed in [11]. To begin with, we introduce some notations and results given in [11].
Definition C.1 (Empirical Rademacher complexity, p. 30 in [11]).
Let $H$ be a family of functions mapping from $\mathcal{Z}$ to $[a,b]$ and $S=(z_{1},...,z_{N})$ a fixed sample of size $N$ with elements in $\mathcal{Z}$. Then, the empirical Rademacher complexity of $H$ with respect to the sample $S$ is defined as:
^$\widehat{\mathcal{R}}_{S}(H)=\mathbb{E}_{\sigma}\left[\sup_{h\in H}\frac{1}{N}\sum_{i=1}^{N}\sigma_{i}h(z_{i})\right],$ (15)
where $\sigma=(\sigma_{1},...,\sigma_{N})$, with $\sigma_{i}$’s independent uniform random variables taking values in $\{-1,+1\}$. The random variables $\sigma_{i}$ are called Rademacher variables.
Theorem 3 (Theorem 3.3, p. 31 in [11]).
Let $\mathcal{G}$ be a family of functions mapping from $\mathbb{Z}$ to $[0,1]$. Then, for any $\delta>0$, with probability at least $1-\delta$ over the draw of an i.i.d. sample $S$ of size $N$, the following holds for all $g\in\mathcal{G}$:
$\mathbb{E}[g(Z)]\leq\frac{1}{N}\sum_{i=1}^{N}g(z_{i})+2\widehat{\mathcal{R}}_{S}(\mathcal{G})+3\sqrt{\frac{\log\frac{2}{\delta}}{2N}}\ .$
Lemma 1 (Talagrand’s lemma, p.93 in [11]).
Let $\Phi:\mathbb{R}\mapsto\mathbb{R}$ be an $l$-Lipschitz ($l>0$). Then, for any hypothesis set $H$ of real-valued functions, the following inequality holds:
^$\widehat{\mathcal{R}}_{S}(\Phi\circ H)\leq l\widehat{\mathcal{R}}_{S}(H)\ .$
Theorem 4 (Margin bound for binary classification, p.94 in [11]).
Let $H$ be a set of real-valued functions and let $P_{X}$ be the distribution over the input space $\mathcal{X}$. Fix $\rho>0$, then, for any $\delta>0$, with probability at least $1-\delta$ over a sample $S$ of size $N$ drawn according to $P_{X}$, the following holds for all $h\in H$:
$R(h)\leq\widehat{R}_{S,\rho}(h)+\frac{2}{\rho}\widehat{\mathcal{R}}_{S}(H)+3\sqrt{\frac{\log\frac{2}{\delta}}{2N}}\ ,$ (16)
where $R(h)=\mathbb{E}[1_{Yh(X)<0}]$ (generalization error), $\widehat{R}_{S,\rho}(h)=\frac{1}{N}\sum_{i=1}^{N}\Phi_{\rho}(y_{i}h(x_{i}))$ is the empirical margin loss on $S$ of size $N$ and ^$\widehat{\mathcal{R}}_{S}(H)$ is the empirical Rademacher complexity of $H$ with respect to the sample $S$
Using Theorem 3 and Lemma 1, we can derive a theorem similar to Theorem 4 but without label as follow:
Theorem 5.
Let $H$ be a set of real-valued functions and let $P_{X}$ be the distribution over the input space $\mathcal{X}$. Fix $\rho>0$, then, for any $\delta>0$, with probability at least $1-\delta$ over a sample $S$ of size $N$ drawn according to $P_{X}$, the following holds for all $h\in H$:
$R(h)\leq\widehat{R}_{S,\rho}(h)+\frac{2}{\rho}\widehat{\mathcal{R}}_{S}(H)+3\sqrt{\frac{\log\frac{2}{\delta}}{2N}}\ ,$ (17)
where $R(h)=\mathbb{E}[1_{\{h(X)<0\}}]$ (generalization error), $\widehat{R}_{S,\rho}(h)=\frac{1}{N}\sum_{i=1}^{N}\Phi_{\rho}(h(x_{i}))$ is the empirical margin loss on $S$ of size $N$ and ^$\widehat{\mathcal{R}}_{S}(H)$ is the empirical Rademacher complexity of $H$ with respect to the sample $S$
Proof.
Let $\mathcal{G}=\{\Phi_{\rho}\circ h:h\in H\}$. By theorem 3, for all $g\in\mathcal{G}$,
$\mathbb{E}[g(X)]\leq\frac{1}{N}\sum_{i=1}^{N}g(x_{i})+2\widehat{\mathcal{R}}_{S}(\mathcal{G})+3\sqrt{\frac{\log\frac{2}{\delta}}{2N}}\ ,$
and consequently, for all $h\in H$,
$\mathbb{E}[\Phi_{\rho}(h(X))]\leq\frac{1}{N}\sum_{i=1}^{N}\Phi_{\rho}(h(x_{i}))+2\widehat{\mathcal{R}}_{S}(\Phi_{\rho}\circ H)+3\sqrt{\frac{\log\frac{2}{\delta}}{2N}}\ .$
Since $1_{\{u\leq 0\}}\leq\Phi_{\rho}(u)$, we have $R(h)=\mathbb{E}[1_{\{h(X)<0\}}]\leq\mathbb{E}[\Phi_{\rho}(h(X))]$, so
$R(h)\leq\widehat{R}_{S,\rho}(h)+2\widehat{\mathcal{R}}_{S}(\Phi_{\rho}\circ H)+3\sqrt{\frac{\log\frac{2}{\delta}}{2N}}\ .$
On the other hand, since $\Phi_{\rho}$ is $1/\rho$-Lipschitz, by lemma 1, we have ^$\widehat{\mathcal{R}}_{S}(\Phi_{\rho}\circ H)\leq\frac{1}{\rho}\widehat{\mathcal{R}}_{S}(H)$. So, $R(h)\leq\widehat{R}_{S,\rho}(h)+\frac{2}{\rho}\widehat{\mathcal{R}}_{S}(H)+3\sqrt{\frac{\log\frac{2}{\delta}}{2N}}\ .$
With all these notions and theorems, we are well-equipped to prove theorems 1 and 2.
C.1 Proof of theorem 1
By using Theorem 4, to prove Theorem 1, it suffices to prove the following result:
Theorem 6.
Each of the following hold:
1. Case where $f\in\mathcal{G}_{1}$ is fixed. The empirical Rademacher complexity of $H_{f}$ can be bounded as follows:
^$\widehat{\mathcal{R}}_{S}(H_{f})\leq\frac{\Gamma\sqrt{(r+R)^{2}+1}}{\sqrt{N}}\ .$
2. Case where $f\in\mathcal{G}_{1}$ is not fixed. The empirical Rademacher complexity of $H_{1}$ can be bounded as follows:
^$\widehat{\mathcal{R}}_{S}(H_{1})\leq\Gamma\sqrt{(r+R)^{2}+1}\ .$
Proof.
Let $\sigma=(\sigma_{1},...,\sigma_{N})$, with $\sigma_{i}$’s independent uniform random variables taking values in $\{-1,+1\}$. Now we calculate the Rademacher complexity.
1. Case where   $f\in\mathcal{G}_{1}$   is fixed.
^$\begin{split}\widehat{\mathcal{R}}_{S}(H_{f})&=\mathbb{E}_{\sigma}\left[\sup_{||w^{\prime}||\leq\Gamma}\frac{1}{N}\sum_{i=1}^{N}\sigma_{i}(\langle f(x_{i}),w\rangle+b)\right](\mbox{ where }w^{\prime}:=(w,b))\\ &=\mathbb{E}_{\sigma}\left[\sup_{||w^{\prime}||\leq\Gamma}\frac{1}{N}\sum_{i=1}^{N}\sigma_{i}(\langle(f(x_{i}),1),(w,b)\rangle)\right]=\mathbb{E}_{\sigma}\left[\sup_{||w^{\prime}||\leq\Gamma}\frac{1}{N}\sum_{i=1}^{N}\sigma_{i}(\langle(f(x_{i}),1),w^{\prime}\rangle)\right]\\ &=\frac{1}{N}\mathbb{E}_{\sigma}\left[\sup_{||w^{\prime}||\leq\Gamma}\langle w^{\prime},\sum_{i=1}^{N}\sigma_{i}(f(x_{i}),1)\rangle\right]\leq\frac{1}{N}\mathbb{E}_{\sigma}\left[\sup_{||w^{\prime}||\leq\Gamma}||w^{\prime}||\times\left\|\sum_{i=1}^{N}\sigma_{i}(f(x_{i}),1)\right\|\right]\\ &\leq\frac{\Gamma}{N}\mathbb{E}_{\sigma}\left[\left\|\sum_{i=1}^{N}\sigma_{i}(f(x_{i}),1)\right\|\right]\leq\frac{\Gamma}{N}\left(\mathbb{E}_{\sigma}\left[\left\|\sum_{i=1}^{N}\sigma_{i}(f(x_{i}),1)\right\|^{2}\right]\right)^{\frac{1}{2}}\;\;\;\;\text{(Jensen's inequality)}\\ &=\frac{\Gamma}{N}\left(\mathbb{E}_{\sigma}\left[\sum_{i,j=1}^{N}\sigma_{i}\sigma_{j}\left\langle(f(x_{i}),1),(f(x_{j}),1)\right\rangle\right]\right)^{\frac{1}{2}}\\ &=\frac{\Gamma}{N}\left(\mathbb{E}_{\sigma}\left[\sum_{i=1}^{N}||(f(x_{i}),1)||^{2}\right]\right)^{\frac{1}{2}}\;\;\;\;(\mathbb{E}_{\sigma}[\sigma_{i}\sigma_{j}]=0\ \text{if}\ i\neq j\ \text{and}\ 1\ \text{otherwise})\\ &=\frac{\Gamma}{N}\left(\mathbb{E}_{\sigma}\left[\sum_{i=1}^{N}(||(f(x_{i})||^{2}+1)\right]\right)^{\frac{1}{2}}=\frac{\Gamma}{N}\left(N+\mathbb{E}_{\sigma}\left[\sum_{i=1}^{N}||(f(x_{i})||^{2}\right]\right)^{\frac{1}{2}}\\ \end{split}$
Now, notice that by triangle inequality, we have $||f(x_{i})||\leq||f(x_{i})-m(q(i))||+||m(q(i))||$, where $m(q(i))\in\{m_{1},m_{2}\}$ is the center of the hypersphere of radius $r$ containing $q_{i}$ (recall that $q_{i}=f(x_{i})$). This means that $||f(x_{i})-m(q(i))||\leq r$ and $||m(q(i))||\leq R$. So, $||f(x_{i})||\leq R+r$. Thus, ^$\widehat{\mathcal{R}}_{S}(H_{f})\leq\frac{\Gamma}{N}\left(N+\mathbb{E}_{\sigma}[\sum_{i=1}^{N}(R+r)^{2}]\right)^{\frac{1}{2}}=\frac{\Gamma}{N}\sqrt{N+N(R+r)^{2}}=\frac{\Gamma\sqrt{(r+R)^{2}+1}}{\sqrt{N}}\ .$
2. Case where   $f\in\mathcal{G}_{1}$   is not fixed.
^$\begin{split}\widehat{\mathcal{R}}_{S}(H)&=\mathbb{E}_{\sigma}\left[\sup_{||w^{\prime}||\leq\Gamma,f\in\mathcal{G}_{1}}\frac{1}{N}\sum_{i=1}^{N}\sigma_{i}(\langle f(x_{i}),w\rangle+b)\right]=\frac{1}{N}\mathbb{E}_{\sigma}\left[\sup_{||w^{\prime}||\leq\Gamma,f\in\mathcal{G}_{1}}\sum_{i=1}^{N}\sigma_{i}(\langle(f(x_{i}),1),w^{\prime}\rangle)\right]\\ &=\frac{1}{N}\mathbb{E}_{\sigma}\left[\sup_{||w^{\prime}||\leq\Gamma,f\in\mathcal{G}_{1}}\langle w,\sum_{i=1}^{N}\sigma_{i}(f(x_{i}),1)\rangle\right]\leq\frac{1}{N}\mathbb{E}_{\sigma}\left[\sup_{||w^{\prime}||\leq\Gamma,f\in\mathcal{G}_{1}}||w^{\prime}||\times\left\|\sum_{i=1}^{N}\sigma_{i}(f(x_{i}),1)\right\|\right]\\ &\leq\frac{\Gamma}{N}\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{G}_{1}}\left\|\sum_{i=1}^{N}\sigma_{i}(f(x_{i}),1)\right\|\right]\leq\frac{\Gamma}{N}\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{G}_{1}}\sum_{i=1}^{N}||\sigma_{i}(f(x_{i}),1)||\right]=\frac{\Gamma}{N}\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{G}_{1}}\sum_{i=1}^{N}||(f(x_{i}),1)||\right]\\ &\leq\frac{\Gamma}{N}\mathbb{E}_{\sigma}\left[\sum_{i=1}^{N}\sup_{f\in\mathcal{G}_{1}}||(f(x_{i}),1)||\right]=\frac{\Gamma}{N}\mathbb{E}_{\sigma}\left[\sum_{i=1}^{N}\sup_{f\in\mathcal{G}_{1}}\sqrt{||(f(x_{i})||^{2}+1}\right]\\ &\leq\frac{\Gamma}{N}\mathbb{E}_{\sigma}\left[\sum_{i=1}^{N}\sqrt{(r+R)^{2}+1}\right]\;\;\;\;(\sup_{f\in\mathcal{G}_{1}}||f(x_{i})||\leq r+R\,\ \text{similarly to {Case 1}})\\ &=\frac{\Gamma}{N}\times N\sqrt{(r+R)^{2}+1}=\Gamma\sqrt{(r+R)^{2}+1}\end{split}$
C.2 Proof of Theorem 2
By using Theorem 5, to prove Theorem 2, it suffices to prove the following result:
Theorem 7.
The empirical Rademacher complexity of $H_{2}$ can be bounded as follows:
^$\widehat{\mathcal{R}}_{S}(H_{2})\leq\Lambda^{2}+2R\Lambda+\frac{R^{2}}{\sqrt{N}}\ .$ (18)
Proof.
Let $\mathcal{G}_{2}=\{f:||\sup_{x\in\mathcal{X}}f(x)||\leq\Lambda\}$. Let $\sigma=(\sigma_{1},...,\sigma_{N})$, with $\sigma_{i}$’s independent uniform random variables taking values in $\{-1,+1\}$. By definition of the empirical Rademacher complexity, we have:
^$\begin{split}\widehat{\mathcal{R}}_{S}(H_{2})&=\mathbb{E}_{\sigma}\left[\sup_{||m||\leq R,f\in\mathcal{G}_{2}}\frac{1}{N}\sum_{i=1}^{N}\sigma_{i}(r^{2}-||f(x_{i})-m||^{2})\right]\\ &=\frac{1}{N}\mathbb{E}_{\sigma}\left[\sum_{i=1}^{N}\sigma_{i}r^{2}+\sup_{||m||\leq R,f\in\mathcal{G}_{2}}\sum_{i=1}^{N}-\sigma_{i}||f(x_{i})-m||^{2}\right]\\ &=\frac{1}{N}\mathbb{E}_{\sigma}\left[\sup_{||m||\leq R,f\in\mathcal{G}_{2}}\sum_{i=1}^{N}-\sigma_{i}||f(x_{i})-m||^{2}\right]\;\;\;\;\left(\mbox{ as }\mathbb{E}_{\sigma}\left[\sum_{i=1}^{N}\sigma_{i}r^{2}\right]=r^{2}\sum_{i=1}^{N}\mathbb{E}_{\sigma}[\sigma_{i}]=0\right)\\ &=\frac{1}{N}\mathbb{E}_{\sigma}\left[\sup_{||m||\leq R,f\in\mathcal{G}_{2}}\sum_{i=1}^{N}(-\sigma_{i}||f(x_{i})||^{2}-\sigma_{i}||m||^{2}+2\sigma_{i}\langle f(x_{i}),m\rangle)\right]\\ &\leq\frac{1}{N}\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{G}_{2}}\sum_{i=1}^{N}-\sigma_{i}||f(x_{i})||^{2}+\sup_{||m||\leq R}\sum_{i=1}^{N}-\sigma_{i}||m||^{2}+\sup_{||m||\leq R,f\in\mathcal{G}_{2}}\sum_{i=1}^{N}2\sigma_{i}\langle f(x_{i}),m\rangle\textperiodcentered\right]\end{split}$
Considering each term inside the expectation operation, we get,
$\begin{split}\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{G}_{2}}\sum_{i=1}^{N}-\sigma_{i}||f(x_{i})||^{2}\right]&\leq\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{G}_{2}}|\sum_{i=1}^{N}-\sigma_{i}||f(x_{i})||^{2}|\right]\leq\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{G}_{2}}\sum_{i=1}^{N}|\sigma_{i}||f(x_{i})||^{2}|\right]\\ &=\mathbb{E}_{\sigma}\left[\sup_{f\in\mathcal{G}_{2}}\sum_{i=1}^{N}||f(x_{i})||^{2}\right]\leq\mathbb{E}_{\sigma}\left[\sum_{i=1}^{N}\Lambda^{2}\right]=N\Lambda^{2}\ .\end{split}$
Consider now the second term. We have,
$\begin{split}\mathbb{E}_{\sigma}\left[\sup_{||m||\leq R}\sum_{i=1}^{N}-\sigma_{i}||m||^{2}\right]&\leq\mathbb{E}_{\sigma}\left[\sup_{||m||\leq R}\left|\sum_{i=1}^{N}-\sigma_{i}||m||^{2}\right|\right]=\mathbb{E}_{\sigma}\left[\sup_{||m||\leq R}||m||^{2}\left|\sum_{i=1}^{N}\sigma_{i}\right|\right]\\ &\leq R^{2}\mathbb{E}_{\sigma}\left[\left|\sum_{i=1}^{N}\sigma_{i}\right|\right]\leq R^{2}\left(\mathbb{E}_{\sigma}\left[\left|\sum_{i=1}^{N}\sigma_{i}\right|^{2}\right]\right)^{\frac{1}{2}}=R^{2}\left(\mathbb{E}_{\sigma}\left[\sum_{i,j=1}^{N}\sigma_{i}\sigma_{j}\right]\right)^{\frac{1}{2}}\\ &=R^{2}\left(\mathbb{E}_{\sigma}\left[\sum_{i=1}^{N}\sigma_{i}^{2}\right]\right)^{\frac{1}{2}}\;\;\;\;(\mbox{ as }\mathbb{E}_{\sigma}[\sigma_{i}\sigma_{j}]=0\ \text{if}\ i\neq j\ \text{and}\ 1\ \text{otherwise})\\ &=R^{2}\sqrt{N}\ .\end{split}$
Consider now the final term inside the expectation operation.
$\begin{split}\mathbb{E}_{\sigma}\left[\sup_{||m||\leq R,f\in\mathcal{G}_{2}}\sum_{i=1}^{N}2\sigma_{i}\langle f(x_{i}),m\rangle\right]&\leq\mathbb{E}_{\sigma}\left[\sup_{||m||\leq R,f\in\mathcal{G}_{2}}\left|\sum_{i=1}^{N}2\sigma_{i}\langle f(x_{i}),m\rangle\right|\right]\\ &\leq\mathbb{E}_{\sigma}\left[\sup_{||m||\leq R,f\in\mathcal{G}_{2}}\left|\sum_{i=1}^{N}2\sigma_{i}||f(x_{i})||\cdot||m||\right|\right]\\ &\leq\mathbb{E}_{\sigma}\left[2R\sup_{f\in\mathcal{G}_{2}}\sum_{i=1}^{N}|\sigma_{i}|\cdot||f(x_{i})||\right]\\ &=\mathbb{E}_{\sigma}\left[2R\sup_{f\in\mathcal{G}_{2}}\sum_{i=1}^{N}||f(x_{i})||\right]\leq\mathbb{E}_{\sigma}\left[2R\sum_{i=1}^{N}\Lambda\right]=2NR\Lambda\ .\end{split}$
Hence, ^$\widehat{\mathcal{R}}_{S}(H_{2})\leq\frac{1}{N}(\Lambda^{2}N+R^{2}\sqrt{N}+2NR\Lambda)=\Lambda^{2}+2R\Lambda+\frac{R^{2}}{\sqrt{N}}\ .$